OpenStudy (anonymous):

Is my answer correct? Yes or no? What is the range of the graph of y = 3(x - 3)2 + 4? y _<_ 4

jimthompson5910 (jim_thompson5910):

close, very close, but it should be \[\Large y \ge 4\]

OpenStudy (anonymous):

What do the following two equations represent? 4x−5y=−3 −20x+25y=15 perpendicular parallel none of the above equivalent